how do you solve 7^(x+4)=1?

Answer by stanbon(75887)

how do you solve 7^(x+4)=1?
---
Take the log of both sides:
(x+4)*log(7) = 0
---
x+4 = 0
---
x = -4
===============================
Cheers,
Stan H.
====================